OpenStudy (anonymous):

please help me understand this. 5(x+7) -2(x-1)=14-(1-9x). This is what I have so far 7x+33=14-(1-9x)

OpenStudy (anonymous):

You have done 25% of the problem. Now open the parenthesis on the right side by multiplying the terms inside by-1 (change signs in other words).

OpenStudy (anonymous):

5x + 35 - 2x + 1 = 14 - 1 + 9x Did you get this on your first step? This is the distributive property

OpenStudy (anonymous):

Sorry... 5x + 35 - 2x + 2 = 14 - 1 + 9x

OpenStudy (anonymous):

In went across 5x+25 then -2x +1 then combined them which gave me 7x+33. I did know I had to change the signs. When do you know when to do that"?

OpenStudy (anonymous):

5(x + 7) = 5x + 35 -2(x - 1) = -2x + 2 14 = 14 -(1 - 9x) = -1 + 9x Do you get each of these?

OpenStudy (anonymous):

oops it was 5x +35 sorry

OpenStudy (anonymous):

I',m sorry i got the 5+35 then -2x+2 then I changed the signs and changed the signs and alligned which gave me 7x +33=14-(1-9)

OpenStudy (anonymous):

You can do this one of two ways. -(2x - 2) which is -2x + 2 or -2(x - 1) you take -2(x) and -2(-1) which gives you -2x + 2

OpenStudy (anonymous):

So 5x + 35 - 2x + 2 = 3x + 37 because 5x - 2x = 3x 35 + 2 = 37

OpenStudy (anonymous):

I did it another way 5x +35 -2x+2 ( I changed the signs) +2x-2 which gave me 7x+33

OpenStudy (anonymous):

Why would you change the signs? You already used the - when you multiplied.

OpenStudy (anonymous):

I don't know??? okay it's official Im lost. I know the answer is 4, but I've tried several ways to figure it out

OpenStudy (anonymous):

3x + 37 = 14 - 1 + 9x 3x + 37 = 13 + 9x subtract 3x from both sides 37 = 13 + 6x subtract 13 from both sides 24 = 6x divide by 6 on both sides 4 = x

OpenStudy (anonymous):

I see it now!!!!! Thanks for all of your help...